A cylinder contains 4.0 g of nitrogen gas. A piston compresses the gas to half its initial volume. Afterward, by what factor has the mass density of the gas increased ?
Answer #1

density p1=m/V1 ... (1)

mass and moles are constant since no gas scaped

volume has changed so mass density also changes

V2=V1/2

p2= m/V2 = m/(V1/2) so p2= 2 * m/V1

from (1) p2= 2*p1

Increased by a factor of 2